# Predictive Traffic Rule Compliance using Reinforcement Learning

### v1

@@ Abstract
Autonomous vehicle path planning has reached a stage where safety and regulatory compliance are crucial. This paper presents a new approach that integrates a motion planner with a deep reinforcement learning model to predict potential traffic rule violations. In this setup, the predictions of the critic directly affect the cost function of the motion planner, guiding the choices of the trajectory. We incorporate key interstate rules from the German Road Traffic Regulation into a rule book and use a graph-based state representation to handle complex traffic information. Our main innovation is replacing the standard actor network in an actor-critic setup with a motion planning module, which ensures both predictable trajectory generation and prevention of long-term rule violations. Experiments on an open German highway dataset show that the model can predict and prevent traffic rule violations beyond the planning horizon, significantly increasing safety in challenging traffic conditions.
@@ I Introduction
Recent research has effectively applied temporal logic to formalize traffic rules, enabling automated online monitoring systems [11, 10, 14] to continuously monitor the compliance of traffic rules. This approach uses the concept of rule robustness—a quantitative metric indicating how thoroughly specific traffic rules are satisfied or violated. Integrating such monitors into autonomous driving frameworks can help prevent rule violations within the planning horizon.
However, a significant challenge emerges from the limited prediction horizons of autonomous driving modules, particularly in complex traffic scenarios. Current planning strategies, which respond solely to immediate violations, are unable to anticipate potential infringements beyond their present scope. This limitation underscores the critical need to extend rule compliance prediction beyond the immediate horizon, especially for rules involving multi-object interactions and highly susceptible to abrupt environmental changes.
Building on steady progress in machine learning, there has been remarkable success in various autonomous driving applications, particularly in the interpretation of intricate traffic patterns and prediction of vehicle behaviors [7]. In addition, reinforcement learning (RL) is especially suited for autonomous driving tasks due to its ability to adapt to complex environments, continuously learn and improve, and integrate rules with learning.
Building on these foundations, we propose a new actor-critic-style deep reinforcement learning (DRL) framework that incorporates a hierarchically structured rule book comprising German interstate traffic rules. To facilitate predictive traffic rule compliance, we use the temporal logic robustness of these rules as reward signals, enabling the critic network to learn state-value functions that indicate potential rule violations. The learned value function then guides the actor’s exploration.
Despite the benefits of DRL, agents often struggle with convergence in complex motion planning tasks involving multiple constraints. Although certain methods aim to enhance RL training efficiency [15], the outputs of actor networks can remain unstable, which requires extensive training for satisfactory performance.
To overcome these limitations, we propose a hybrid approach: instead of depending exclusively on DRL for exploration, we incorporate a motion planner as the actor component. In this thesis, we adopt the Commonroad Reactive Planner [21], which has demonstrated both efficiency and reliability in generating robust, explainable trajectories. This planner leverages a lattice-based methodology [20, 19] to produce candidate trajectories and identifies the optimal path via a cost function.
In a conventional actor-critic (AC) approach, critic evaluation provides feedback to the actor (policy network), guiding policy improvements. To integrate an actor-critic framework into our system, we embed the critic’s output directly into the motion planner’s cost function, thus selecting trajectories based on predicted state values. In other words, the planner prefers routes with higher state values, which are more likely to avoid long-term traffic rule violations. This method effectively mirrors on-policy learning by synchronizing the actor’s decisions with updates to the critic network during exploration.
@@
Recognizing the importance of high-quality representations of environmental states in reinforcement learning, we employ a Graph Neural Network (GNN) architecture [13] as a feature extractor. This framework captures arbitrary road layouts and varying numbers of surrounding vehicles, while allowing flexible choices of input features. As a result, our agent benefits from a comprehensive state representation. An overview of the proposed approach is presented in Figure 1.
@@
In this paper, we present a new approach for predicting future traffic rule violations by integrating reinforcement learning with lattice-based motion planning. Our work makes several key contributions to the field of autonomous driving:
@@
    We design a plug-and-play modular predictive traffic rule compliance system that can be integrated with an existing autonomous driving motion planner.
@@
    We pioneer the integration of actor-critic reinforcement learning with planning methods, creating a framework that combines the strengths of both.
@@
To the best of our knowledge, this is the first successful integration of actor-critic reinforcement learning algorithms with cost-based motion planning in autonomous driving. Our work opens new avenues for uniting learning-based prediction with explainable planning methods, particularly for forecasting and preventing future traffic rule violations.
@@ II-A Reinforcement Learning
Reinforcement learning (RL) is a powerful machine learning paradigm that enables an agent to learn optimal decision-making strategies by interacting with an environment to maximize a cumulative reward signal. Unlike supervised learning, which depends on large, prelabeled datasets, RL adopts a trial-and-error approach, allowing the agent to mirror natural learning processes through direct environmental feedback. This framework is particularly well suited for solving Markov Decision Processes (MDPs), which formalize the environment in terms of states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, transition probabilities $P(s^{\prime}|s,a)$, and rewards $r(s,a,s^{\prime})$. The agent’s goal is to derive a policy - a mapping from states to actions $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that optimizes the expected long-term reward [17].
To evaluate the quality of states and pairs of state actions, RL algorithms define value functions. The state value function $V^{\pi}(s)$ represents the expected return of the state $s$ following policy $\pi$:
$V^{\pi}(s)=\mathbb{E}_{\pi}[G_{t}|S_{t}=s]$ (1)
One way to solve MDPs with RL is the actor-critic method. It consists of two complementary components:
  • Actor: Develops a policy $\pi(s)$ that generates actions, typically using a stochastic policy with a distributional output layer.
  • Critic: Evaluates the quality of states through a value function $V(s)$, where $s$ denotes the system state [17].
In standard implementations, the actor approximates the policy $\pi_{\phi}(a|s)$ and the critic approximates the state-value function $V_{\xi}(s)$, both using multilayer perceptrons (MLPs). These components maintain separate training processes but are updated simultaneously. The critic typically learns state values through methods like temporal difference (TD) learning or Monte Carlo estimation, providing feedback to refine the actor’s policy.
@@ II-C Traffic Rules Formalization
Formalizing traffic rules into quantifiable metrics poses a significant hurdle for RL agents. Simple rules, such as speed limits, can be encoded directly through violation penalties, for example, a negative reward for exceeding a threshold. However, more sophisticated rules, such as ”the ego vehicle must maintain sufficient speed to avoid impeding traffic flow,” require comprehensive contextual information and cannot be adequately captured by binary compliance (satisfied or violated). Real-world driving often involves continuous degrees of adherence to the rules, and Boolean representations do not account for reasonable deviations, such as temporarily exceeding a speed limit to avoid an obstacle.
To overcome these limitations, formal temporal logic frameworks have emerged as powerful tools for specifying and evaluating traffic rules. Linear Temporal Logic (LTL) has been used to formulate traffic rules in an automatically evaluable syntax [16, 3]. However, its discrete-time framework limits its suitability for online evaluation in RL, where real-time adaptability is essential. Signal Temporal Logic (STL), as an extension of Metric Temporal Logic (MTL) [8], addresses these shortcomings by specifying properties of continuous-time signals [12]. Unlike MTL’s Boolean satisfaction values, STL introduces quantitative semantics through the degree of robustness, which measures the degree of satisfaction or violation of the rule [5]. This makes STL ideal for autonomous driving, where compliance often exists on a spectrum.
@@ III Methdology
In this section, we introduce our proposed method for predicting and preventing future traffic rule violations in autonomous driving scenarios. Our approach integrates a modified RL framework with a motion planner, leveraging the complementary strengths of both techniques to ensure that traffic rule compliance extends beyond the immediate planning horizon.
Our method employs a Graph Neural Network (GNN) extractor within an actor-critic RL framework, coupled with a motion planner, to train an ego agent to navigate highway scenarios while adhering to hierarchical traffic rules. Specifically, we focus on trucks operating on German interstate highways, where scenarios involve multiple lanes and traffic participants.
Unlike many existing approaches that rely on simulators to model the motion of traffic participants, we utilize a series of recorded scenarios for training instead of interactive scenarios. This choice is intentional: In reactive simulations, other vehicles actively avoid the ego vehicle, potentially simplifying the learning task and leading to overly aggressive or less robust agent behavior. In contrast, recorded scenarios present static, unresponsive traffic participants, creating more challenging conditions that rigorously test the ego agent’s planning and obstacle avoidance capabilities.
The primary objective of our algorithm is to enable the ego vehicle to achieve multiple goals. First, the agent derives a state value from the input features of the current scenario, obtained through sensors and other data sources, that represents the long-term cumulative reward associated with compliance with specific traffic rules. Then, by selecting the trajectory with the highest state value, the ego vehicle leverages the motion planner to generate a path that predictively prevents long-term rule violations.
To enhance this process, we integrate traffic rules with varying weights based on their hierarchical importance, as defined by a rule book. In addition, we incorporate cost functions to ensure that planned trajectories are goal-oriented, safe, and comfortable. Through this comprehensive approach, our goal is to develop a robust planner that delivers exceptional performance in complex driving environments.
@@ III-A Selection and Formulation of Traffic Rules
This study centers on autonomous driving scenarios along the German interstate system, extending the foundational research presented in  [11]. In that work, seven interstate traffic rules were formalized using Signal Temporal Logic (STL), drawing from the German Road Traffic Regulation (StVO), the Vienna Convention on Road Traffic (VCoRT), and legal expertise. These rules address critical aspects of driving, including safety, speed limits, traffic flow, and local customs. However, not all of these formalized rules pose significant challenges or are equally relevant to our proposed methodology.
@@
Consequently, our approach emphasizes scenarios where conventional rule-encoding strategies are inadequate, highlighting the advantages of our data-driven, neural network-based methodology. For example, a vehicle with a 2 seconds prediction and planning horizon cannot effectively handle a rule violation at 2.1 seconds, since its motion planning and control systems typically cannot adjust in time, inevitably leading to a rule violation. To this end, we have carefully selected two representative rules from  [11] that pose more complex challenges, requiring a sophisticated integration of traffic regulations into the autonomous vehicle decision-making process. Furthermore, we introduce a newly formalized rule to underscore the critical role of our method in addressing intricate traffic scenarios. This selection and formulation highlight the shortcomings of traditional approaches and illustrate the potential of our advanced techniques to improve autonomous driving performance on German interstates.
@@
The detailed formulas for these rules, formalized using Signal Temporal Logic (STL), are presented in Table I.
@@ III-B Graph Representation and GNN
In this section, we introduce the method used to embed traffic scenario input features into a meaningful representation using Graph Neural Networks (GNNs).
To begin, we construct a graph representation of the traffic scenario. A graph consists of nodes and edges. In the context of traffic, the nodes could represent vehicles, lanes, or other participants such as bicycles and pedestrians. However, given the location on the highway, we exclude non-vehicle participants. Furthermore, the relatively simple layout of the lane of the roads, mainly consisting of roads and ramps, makes the use of lanelets unnecessary. Representing lanelets as nodes would introduce multiple node types, resulting in a heterogeneous graph, which is more complex to process with GNNs. Consequently, we simplify the representation by using only vehicle nodes connected by vehicle-to-vehicle (V2V) edges to form the graph.
In the work by  [6], the $V2V$ graph is constructed using basic kinematic vehicle features and relative position edge features, as detailed in Table II. However, these features alone are insufficient for the model to develop a comprehensive understanding required to predict the compliance with traffic rules. To address this limitation, we enhance the graph representation by incorporating additional features designed to help the model recognize specific patterns associated with rule compliance.
Drawing inspiration from  [15], who employ rule robustness as node features and predicates such as CutIn and InFrontOf as edge features, we initially considered a similar approach. However, the computational overhead of repeatedly extracting features and calculating robustness proved prohibitive for our method, which involves numerous iterative operations. Instead, we adopt simpler, yet targeted, features tailored to our specific traffic rules. The enhanced graph representation, including these additional features, is also presented in Table II.
@@
Intuitively, the ego vehicle has access to more detailed information than other agents. Therefore, in addition to standard node and edge features, we incorporate extensive ego-specific features, which are later combined with the embedded features from the GNN. Key additions include features such as DistLeftBound, DistRightBound, DistLeftRoadBound, and DistRightRoadBound, which are added to both vehicle and ego features to enable vehicles to localize themselves within the lanelet network. Since lanelets are not represented as nodes, these features are essential for recognizing lane-related rules, all of which are relevant to our selected rules. Additionally, a Lane Feature explicitly indicates the lane occupied by each vehicle, facilitating the assessment of rule compliance. For traffic rule $R_{I6}$, we introduce the NonOvertakingTSRelative feature, which provides the relative longitudinal distance to the next “no overtaking” traffic sign, based on a sensor radius of 100 meters. This range ensures the vehicle has sufficient time to maneuver and avoid violations. The distance resets to the next relevant sign upon encountering a “no overtaking end” sign. Furthermore, HeadingError and goal-related features enhance the ego vehicle’s situational awareness and support goal-oriented planning.
For $V2V$ edge features, we introduce the RelativeVelocity feature, which is critical for ensuring the ego vehicle does not exceed the speed of the vehicle to its left, supporting compliance with rule $R_{I2}$. The LeftOf feature mimics the predicate calculation for “LeftOf,” providing expressive information with low computational complexity. Additionally, the SameLane feature aids in understanding rule $R\_G2$, which relates to interactions with vehicles in the same lane.
@@ III-C Modified Actor-Critic Algorithm
The actor-critic method is a widely recognized approach in deep reinforcement learning (DRL). In this framework, the critic learns a value function that guides the actor’s policy optimization. During training, these two networks are deeply interdependent: the critic’s value estimates shape the actor’s policy updates, while the actor’s actions generate new experiences that refine the critic’s value assessments. This mutual learning process typically drives both networks toward optimal behavior.
However, applying this method to autonomous driving presents significant challenges. When a neural network serves as the actor, its black-box nature raises concerns about interpretability and controllability. Integrating prior knowledge, such as traffic rules, becomes difficult, potentially leading to unsafe or unstable vehicle behaviors that are difficult to debug or constrain. Furthermore, a randomly initialized neural network actor begins by making random decisions, which impedes effective exploration. This issue is particularly pronounced in interstate traffic scenarios, where the navigable space is narrow relative to the length of the route. Random actions often result in off-road terminations, preventing the agent from reaching its goal.
@@
The effectiveness of this replacement stems from the fact that the output of the value network approximates the state value, which represents the cumulative rewards following the policy. When using robustness of traffic rules as a reward, a higher state value indicates that the vehicle is more likely to comply with rules in that state, while a lower value suggests the opposite. This information is essential to avoid long-term violations of traffic rules. Consequently, in the training phase of our approach, the motion planner selects trajectories using a state-value-guided mechanism.
Let $V^{*}(s)$ represent the output of the well-trained value function for a state $s$, and let ${\tau_{1},\tau_{2},\dots,\tau_{n}}$ denote the set of candidate trajectories generated by the planner. Each trajectory $\tau_{i}$ is a sequence of collision-free and kinematically feasible states, verified by a previously implemented checker. We define the cost for each trajectory $\tau_{i}$ as follows:
$C(\tau_{i})=\sum_{s_{ij}\in\tau_{i}}(-w_{ij}*V^{*}(s_{ij})),$ (2)
where $w_{ij}$ are the weighting factors assigned to each state. The optimal trajectory is selected as:
$\tau^{*}=\arg\min_{\tau_{i}}C(\tau_{i}).$ (3)
This selection process identifies a macro-action, a trajectory spanning multiple time steps that balances long-term value with prior knowledge. The motion planner constrains the action space to a finite set of trajectories satisfying kinematic and safety requirements. By incorporating $-V^{*}(s_{ij})$ into the cost function, an estimate of future returns is included for each state $s_{ij}$ along the trajectory $\tau_{i}$. The critic then updates its value network based on rollouts collected through the actor’s exploration. This approach aligns with the actor-critic framework’s core principle: the actor selects actions according to the current policy, the critic evaluates the outcomes and estimates their quality, and the actor refines its policy using the critic’s feedback.
During the exploration phase, rather than employing a $\varepsilon$-greedy strategy, a greedy strategy is adopted with random start and goal settings and a planned route. Since the motion planner generates trajectories around the route, randomly created routes ensure comprehensive coverage of all possible states. Furthermore, given the relatively straightforward nature of interstate scenarios, this greedy strategy suffices to encompass nearly all potential situations.
After each exploration step, rule-specific robustness rewards are computed for every state. The training procedure alternates between the exploration and update phases, collecting rollouts prior to updating the network. The critic network is trained to minimize the temporal difference (TD) error.
Due to the real-time updates of both the actor and critic, and since the critic’s state value updates depend on the next states produced by the current policy, this method achieves on-policy learning while maintaining high sample efficiency.
Our method offers significant flexibility. In conventional RL approaches, it is necessary to integrate all rewards that represent multiple planning goals. This often creates challenges when attempting to incorporate new planning goals or constraints, as it typically requires retraining the model from scratch.
@@
In the inference phase, we can flexibly combine the trained value networks. Additionally, we can incorporate supplementary costs related to goal orientation, safety, and efficiency. This integration results in a comprehensive and practical motion planner with enhanced explainability. Leveraging the proposed rule book, the inference model can be constructed with the following cost function:
$\begin{split}C(\tau_{i})&=w_{1}V_{R\_G1}+w_{2}V_{R\_I6}+w_{3}V_{R\_I3}+\sum_{i}w_{i}C_{i},\\ &\quad\text{where}\quad w_{1}>w_{2}>w_{3}\end{split}$ (4)
Here, $V$ represents the output of the trained value function, and $C_{i}$ denotes any additional cost function.

### v2

@@ Abstract
Autonomous vehicle path planning has reached a stage where safety and regulatory compliance are crucial. This paper presents an approach that integrates a motion planner with a deep reinforcement learning model to predict potential traffic rule violations. Our main innovation is replacing the standard actor network in an actor-critic method with a motion planning module, which ensures both stable and interpretable trajectory generation. In this setup, we use traffic rule robustness as the reward to train a reinforcement learning agent’s critic, and the output of the critic is directly used as the cost function of the motion planner, which guides the choices of the trajectory. We incorporate some key interstate rules from the German Road Traffic Regulation into a rule book and use a graph-based state representation to handle complex traffic information. Experiments on an open German highway dataset show that the model can predict and prevent traffic rule violations beyond the planning horizon, increasing safety and rule compliance in challenging traffic scenarios.
@@ I Introduction
Recent research has effectively applied temporal logic to formalize traffic rules, enabling automated online monitoring systems [1, 2, 3] to continuously monitor the compliance of traffic rules. These approaches use the concept of rule robustness—a quantitative metric indicating how thoroughly specific traffic rules are satisfied or violated. Integrating such monitors into autonomous driving frameworks can help prevent rule violations within the planning horizon.
However, a significant challenge emerges from the limited prediction horizons of autonomous driving modules, particularly in complex traffic scenarios. Current planning strategies, which respond solely to immediate violations, are unable to anticipate potential infringements beyond their present scope.
Building on steady progress in machine learning, there has been remarkable success in various autonomous driving applications, particularly in the interpretation of intricate traffic patterns and prediction of vehicle behaviors [4]. Among various machine learning methods, reinforcement learning (RL) is especially suited for autonomous driving tasks due to its ability to adapt to complex environments, continuously learn and improve, and integrate rule information by setting proper rewards.
Building on these foundations, we propose to use an actor-critic (AC) framework that incorporates a hierarchically structured rule book comprising German interstate traffic rules. To facilitate predictive traffic rule compliance, we use the temporal logic robustness of these rules as reward signals, enabling the critic network to learn state-value functions that indicate potential rule violations. The learned value function then guides the actor’s exploration.
Despite the benefits of deep reinforcement learning (DRL), agents often struggle with convergence in complex motion planning tasks involving multiple constraints. Although certain methods aim to enhance RL training efficiency [5], the outputs of actor networks can remain unstable, which requires extensive training for satisfactory performance.
To overcome these limitations, we propose a hybrid approach: we incorporate a cost-based motion planner as the actor component. In this thesis, we demonstrate our idea by adopting the Commonroad Reactive Planner [6], which has demonstrated both efficiency and reliability in generating robust, explainable trajectories. This planner leverages a lattice-based methodology [7, 8] to produce candidate trajectories and identifies the optimal path via a cost function.
@@
Recognizing the importance of high-quality representations of environmental states in RL, we employ a Graph Neural Network (GNN) architecture [9] as a feature extractor. This framework captures arbitrary road layouts and varying numbers of surrounding vehicles, while allowing flexible choices of input features. As a result, our agent benefits from a comprehensive state representation. An overview of the proposed approach is presented in Figure 1.
@@
In this paper, we present a new approach for predicting future traffic rule violations by integrating RL with motion planning for exploration. Our work makes several key contributions to the field of autonomous driving:
@@
    We pioneer the integration of AC with planning methods, creating a framework that combines the strengths of both.
@@
To the best of our knowledge, this is the first successful integration of AC algorithms with cost-based motion planning in autonomous driving. Our work opens new avenues for uniting learning-based prediction with explainable planning methods, particularly for forecasting and preventing future traffic rule violations.
@@ II-A Reinforcement Learning
RL is a powerful machine learning paradigm that enables an agent to learn optimal decision-making strategies by interacting with an environment to maximize a cumulative reward signal. Unlike supervised learning, which depends on large, prelabeled datasets, RL adopts a trial-and-error approach, allowing the agent to mirror natural learning processes through direct environmental feedback. This framework is particularly well suited for solving Markov Decision Processes (MDPs), which formalize the environment in terms of states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, transition probabilities $P(s^{\prime}|s,a)$, and rewards $r(s,a,s^{\prime})$. The agent’s goal is to derive a policy - a mapping from states to actions $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that optimizes the expected long-term reward [10].
To evaluate the quality of states and pairs of state actions, RL algorithms define value functions. The state value function $V^{\pi}(s)$ represents the expected return of the state $s$ following policy $\pi$.
One way to solve MDPs with RL is the AC method. It consists of two complementary components: The actor develops a policy $\pi(s)$ that generates actions, typically using a stochastic policy with a distributional output layer. The critic evaluates the quality of states through a value function $V(s)$, where $s$ denotes the system state [10].
In a standard AC implementation, the actor approximates the policy $\pi_{\phi}(a|s)$ and the critic approximates the state-value function $V_{\xi}(s)$, both using multilayer perceptrons (MLPs). These components maintain separate training processes but are updated simultaneously. The critic typically learns state values through methods like temporal difference (TD) learning or Monte Carlo estimation, providing feedback to refine the actor’s policy.
@@ II-C Traffic Rules Formalization
Formalizing traffic rules into quantifiable metrics poses a significant hurdle for RL agents. Simple rules, such as speed limits, can be encoded directly through violation penalties, for example, a negative reward for exceeding a threshold. However, more sophisticated rules, such as ”the ego vehicle must maintain sufficient speed to avoid impeding traffic flow,” require comprehensive contextual information and cannot be adequately captured by binary compliance. Real-world driving often involves continuous degrees of adherence to the rules, and Boolean representations do not account for reasonable deviations, such as temporarily exceeding a speed limit to avoid an obstacle.
To overcome these limitations, formal temporal logic frameworks have emerged as powerful tools for specifying and evaluating traffic rules. Linear Temporal Logic (LTL) has been used to formulate traffic rules in an automatically evaluable syntax [12, 13]. However, its discrete-time framework limits its suitability for online evaluation in RL. Signal Temporal Logic (STL), as an extension of Metric Temporal Logic (MTL) [14], addresses these shortcomings by specifying properties of continuous-time signals [3]. STL introduces quantitative semantics through the degree of robustness, which measures the degree of satisfaction or violation of the rule [15]. This makes STL ideal for autonomous driving.
@@ III-A Selection and Formulation of Traffic Rules
This study centers on driving scenarios along the German interstate system, extending the foundational research presented in  [1]. In that work, seven interstate traffic rules were formalized using STL, drawing from the German Road Traffic Regulation (StVO), the Vienna Convention on Road Traffic (VCoRT), and legal expertise. These rules address critical aspects of driving, including safety, speed limits, traffic flow, and local customs. However, not all of these formalized rules pose significant challenges or are equally relevant to our proposed methodology.
@@
Consequently, our approach emphasizes scenarios where conventional non-learning strategies are inadequate, highlighting the advantages of our data-driven, neural network-based methodology. For example, a vehicle with a 2 seconds prediction and planning horizon cannot effectively handle a rule violation at 2.1 seconds, since its motion planning and control systems typically cannot adjust in time, inevitably leading to a rule violation. To this end, we have carefully selected two representative rules from  [1] that pose more complex challenges, requiring a sophisticated integration of traffic regulations into the autonomous vehicle decision-making process. Furthermore, we introduce a newly formalized rule to underscore the critical role of our method in addressing intricate traffic scenarios. This selection and formulation highlight the shortcomings of traditional approaches and illustrate the potential of our advanced techniques to improve autonomous driving performance on German interstates.
@@
The detailed formulas for these rules, formalized using STL, are presented in Table I.
@@ III-B Graph Representation and GNN
We construct a graph representation of the traffic scenario. A graph consists of nodes and edges. In the context of traffic, the nodes could represent vehicles, lanes, or other participants such as bicycles and pedestrians. However, given the location on the highway, we exclude non-vehicle participants. Furthermore, the relatively simple layout of the lane of the roads, mainly consisting of roads and ramps, makes the use of lanelets unnecessary. Representing lanelets as nodes would introduce multiple node types, resulting in a heterogeneous graph, which is more complex to process with GNNs. Consequently, we simplify the representation by using only vehicle nodes connected by vehicle-to-vehicle (V2V) edges to form the graph.
In [18], the $V2V$ graph is constructed using basic kinematic vehicle features and relative position edge features, as detailed in Table II. However, these features alone are insufficient for the model to develop a comprehensive understanding required to predict compliance with traffic rules. To address this limitation, we enhance the graph representation by incorporating additional features designed to help the model recognize specific patterns associated with rule compliance.
Drawing inspiration from  [5], which employs rule robustness as node features and predicates such as CutIn and InFrontOf as edge features, we initially considered a similar approach. However, the computational overhead of repeatedly extracting features and calculating robustness proved prohibitive for our method, which involves numerous iterative operations. Instead, we adopt simpler, yet targeted, features tailored to our specific traffic rules. The enhanced graph representation, including these additional features, is also presented in Table II.
@@
Intuitively, the ego vehicle has access to more detailed information than other agents. Therefore, in addition to standard node and edge features, we incorporate extensive ego-specific features, which are later combined with the embedded features from the GNN. Key additions include features such as DistLeftBound, DistRightBound, DistLeftRoadBound, and DistRightRoadBound, which are added to both vehicle and ego features to enable vehicles to localize themselves within the lanelet network. Since lanelets are not represented as nodes, these features are essential for recognizing lane-related rules, all of which are relevant to our selected rules. Additionally, a Lane Feature explicitly indicates the lane occupied by each vehicle, facilitating the assessment of rule compliance. For traffic rule $R\_{I6}$, we introduce the NonOvertakingTSRelative feature, which provides the relative longitudinal distance to the next “no overtaking” traffic sign, based on a sensor radius of 100 meters. This range ensures the vehicle has sufficient time to maneuver and avoid violations. The distance resets to the next relevant sign upon encountering a “no overtaking end” sign. Furthermore, HeadingError and goal-related features enhance the ego vehicle’s situational awareness and support goal-oriented planning.
For $V2V$ edge features, we introduce the RelativeVelocity feature, which is critical for ensuring the ego vehicle does not exceed the speed of the vehicle to its left, supporting compliance with rule $R\_{I2}$. The LeftOf feature mimics the predicate calculation for “LeftOf,” providing expressive information with low computational complexity. Additionally, the SameLane feature aids in understanding rule $R\_G2$, which relates to interactions with vehicles in the same lane.
@@ III-C Modified Actor-Critic Algorithm
Applying AC method to autonomous driving presents significant challenges. When a neural network serves as the actor, its black-box nature raises concerns about interpretability and controllability. Integrating prior knowledge, such as collision avoidance, becomes difficult, leading to unsafe or unstable vehicle behaviors that are impossible to debug or constrain. Furthermore, a randomly initialized neural network actor begins by making random decisions, which impedes effective exploration. This issue is particularly pronounced in interstate traffic scenarios, where the navigable space is narrow relative to the length of the route. Random actions often result in off-road terminations, preventing the agent from reaching its goal.
@@
The effectiveness of this replacement stems from the fact that the output of the value network approximates the state value, which represents the cumulative rewards following the policy. When using robustness of traffic rules as a reward, a higher state value indicates that the vehicle is more likely to comply with rules in that state, while a lower value suggests the opposite, which is essential to avoid long-term violations of traffic rules. Consequently, in the exploration phase of our approach, the motion planner selects trajectories using a state-value-guided mechanism.
This selection process identifies a ”macro-action”, a trajectory spanning multiple time steps that balances long-term value with prior knowledge. The motion planner constrains the action space to a finite set of trajectories satisfying kinematic and safety requirements. By incorporating $-V^{*}(s_{ij})$ into the cost function, an estimate of future returns is included for each state $s_{ij}$ along the trajectory $\tau_{i}$. The critic then updates its value network based on rollouts collected through the actor’s exploration. This approach aligns with the AC framework’s core principle: the actor selects actions according to the current policy, the critic evaluates the outcomes and estimates their quality, and the actor refines its policy using the critic’s feedback.
Moreover, during the exploration phase, rather than employing a $\varepsilon$-greedy strategy, a greedy strategy is adopted with random start and goal settings and a planned route. Since the motion planner generates trajectories around the route, the created routes with random settings ensure comprehensive coverage of possible states.
Due to the simultaneous updates of both the actor and critic, and since the critic’s state value updates depend on the next states produced by the current policy, this method achieves on-policy learning while maintaining high sample efficiency.